***For calculator work, show the entries or steps you use.***

#3. To buy your dream home you use a 4/16 reset loan that pays interest only (bullet loan) for the first 4 years, followed with a reset payment to amortize over the remaining 16 years. The loan amount is $180,000 at 5.5%, with 2 discount points (to the lender) and $875 in third party closing costs which the borrower must pay.

a) What net amount does the lender disburse?

b) What net amount do you as a borrower receive?

c) What is your monthly payment during the first 4 years? d) What is your monthly payment during the remaining 16 years?

e) What is the FTLAPR on this loan?

f) What is your balance 6 years after you take out the loan g) If you repay this loan after 6 years, what is the yield (expressed as an APR) for the lender?

h) What is the yield expressed as an EAR? i) If you repay this loan after 6 years, what is the effective borrower cost (EBC) to you (expressed as an APR)?

j) What is the EBC expressed as an EAR?
Answer #1

a) What net amount does the lender disburse? 0.98*180000 = 176,400
b) What net amount do you as a borrower receive? 176400-875=175,525
c) What is your monthly payment during the first 4 years?
PMT(N=48, I/YR=5.5, PV=-180000, FV=180000) = 825.00
d) What is your monthly payment during the remaining 16 years?
PMT(N=192, I/YR=5.5, PV=-180000, FV=0) = 1411.75
e) What is the FTLAPR on this loan? t CFj Ni 0 -176400.00 1-48 825.00 48 49-144 1411.75 96 145-240 1411.75 96 IRR/YR = 5.72%